For a collected data of 10 random samples , we have 

  0.700 , mean is 0.742, SD is 0.040 . Test the statistic:


1
Expert's answer
2021-01-18T10:13:12-0500

"n = 10 \\

μ = 0.700 \\

\bar{x}=0.742 \\

σ = 0.040 \\

Test-statistic \\

X = \frac{\bar{x}-μ}{σ/ \sqrt{n}} \\

= \frac{0.742-0.700}{0.040/ \sqrt{10}} \\

= \frac{0.042}{0.0126} \\

= 3.334"

Answer: 3.334